OpenStudy (anonymous):

In the following problem, what property allows you to add the 17 and the 3 together before adding the 15? 17 + 15 + 3

OpenStudy (amistre64):

associative i believe; which says that it doesnt matter how you add up a group, it still comes out the same

OpenStudy (bahrom7893):

Theorem : All numbers are equal to zero. Proof: Suppose that a=b. Then a = b a^2 = ab a^2 - b^2 = ab - b^2 (a + b)(a - b) = b(a - b) a + b = b a = 0 Furthermore if a + b = b, and a = b, then b + b = b, and 2b = b, which mean that 2 = 1.

OpenStudy (anonymous):

In the following problem, what property allows you to multiply 5 by 2 before multiplying by 13? 5 • 13 • 2

myininaya (myininaya):

(a+b)+c=a+(b+c) this is associative a+b+c=a+c+b this is commutative

OpenStudy (bahrom7893):

commutativity is the next answer

OpenStudy (anonymous):

its either Associative or commutative

OpenStudy (bahrom7893):

5 • 13 • 2 is commutative property.

OpenStudy (anonymous):

What property allows you to move the parentheses in the following problem? (8 + 7) + 13 = 8 + (7 + 13)

myininaya (myininaya):

look i what i typed above what looks like this?

myininaya (myininaya):

do you see the two rules i typed (and of course there are many more)?